# Ca2-VDM: Efficient Autoregressive Video Diffusion Model with Causal Generation and Cache Sharing

### v1

@@ 3.3 Autoregressive Inference with Cache Sharing
Figure 4: Illustration of causal temporal attention (a) & (b) and the temporal KV-cache queue with Cyclic-TPEs (c). In (c), assume that $L_{\text{train}}=P_{\text{max}}+l$ and $P_{k+l}=P_{k}+l$. We show the state that autoregressive inference reaches $P_{k}=P_{\text{max}}$.
@@
Figure 6: Qualitative examples generated by GenLV [42], StreamT2V [15], OS-Fix, and our Ca2-VDM. We sampled 32 frames with an interval of 8 frames for display. Note that GenLV does not strictly follow the given first frame, since it was not finetuned on explicitly injected conditional frames. In the implementation of GenLV, we used DDIM inversion to build the initial noise based on the first frame.
@@ 4.1 Experimental Setup
Model Details and Baselines. We built Ca2-VDM based on spatial-temporal Transformer following [24, 4] and initialized it with Open-Sora v1.0 [55]. Following PixArt-$\alpha$ [4], we used T5 [31] as the text encoder and used the VAE from StableDiffusion [33]. The length of the clean prefix was randomly sampled according to the multiples of chunk length $l$, i.e., $P\in\{1,1+l,\ldots,1+nl\}$ and $P_{\text{max}}=1+nl$. We used training videos of various lengths with $L_{\text{train}}=P+l$. As comparisons, we built two bidirectional baselines (cf. Figure 1(a)) based on the same Open-Sora v1.0: One was trained with fixed-length conditional frames (denoted as OS-Fix), where $P$ is fixed as $P=L_{\text{train}}/2$ in training and inference. The other was trained with autoregressively extendable conditional frames using the same training configs as Ca2-VDM (denoted as OS-Ext).
Table 1: Zero-shot FVD scores on MSR-VTT [47] and UCF101 [39] test sets. All methods generate video at a resolution of 16$\times$256$\times$256. C: condition. T and I are text and image conditions, respectively.
@@
ModelScope [43] T 550 410
VideoComposer [44] T 580 -
Video-LDM [3] T - 550.6
PYoCo [9] T - 355.2
Make-A-Video [34] T - 367.2
AnimateAnything [6] T+I 443 -
PixelDance [50] T+I 381 242.8
SEINE [5] T+I 181 -
@@
Table 2: Finetuned FVD scores on UCF-101 [39] test set. Methods with were trained on both train and test sets.
@@
MCVD [41] $64^{2}$ 1143
VDT [21] $64^{2}$ 225.7
DIGAN [48] $128^{2}$ 577
TATS [8] $128^{2}$ 420
VideoFusion [23] $128^{2}$ 220
LVDM [14] $256^{2}$ 372
PVDM [49] $256^{2}$ 343.6
Latte [24] $256^{2}$ 333.6
@@
Table 3: Ablations of $P_{\text{max}}$ and prefix-enhancement (PE) on SkyTimelapse [51]. Each variant of Ca2-VDM generated 48 frames with 6 AR steps. The results were divided into three 16-frame chunks for FVD evaluation.
@@
[…]
Evaluation Datasets and Metrics. We used MSR-VTT [47], UCF101 [39], and SkyTimelapse [51] datasets at resolution 256$\times$256, and reported Fréchet Video Distance (FVD) [40] following previous works [50, 9, 5]. More details about choosing text prompts and computing FVD scores on these datasets are left in the Appendix.
4.2 Evaluation for Generation Quality
In-Chunk Generation Quality. We evaluated the zero-shot text-to-video (T2V) FVD scores on MSR-VTT [47] and UCF101 [39], as shown in Table 3. We compared Ca2-VDM to state-of-the-art T2V models including two groups: 1) Text conditioned: ModelScope [43], VideoComposer [44], Video-LDM [3], PYoCO [9], and Make-A-Video [34]. 2) Text with extra image conditioned, e.g., for image-to-video: AnimateAnything [6], PixelDance [50] and video transition: SEINE [5]. We also finetuned Ca2-VDM on UCF101 at resolution 16$\times$256$\times$256 and reported the FVD scores in Table 3. We compared it with SOTA video generation models: MCVD [41], VDT [21], DIGAN [48], TATS [8], LVDM [14], PVDM [49], and Latte [24]. The FVD results in both Table 3 and Table 3 show that our Ca2-VDM has a competitive T2V performance with SOTA models. We further show some qualitative examples in Figure 6. We can see that Ca2-VDM has comparable generation quality with existing SOTA models.
Table 4: FVD results on MSR-VTT [47] test set.
Method FVD between AR step 1 and $i$
$i=2$ $i=3$ $i=4$ $i=5$ $i=6$
GenLV [42] 282.8 291.4 299.0 318.2 310.3
StreamT2V [15] 317.5 434.7 478.2 462.0 512.4
OS-Fix 182.9 210.6 260.8 284.3 315.1
Ca2-VDM 160.6 206.5 262.8 281.3 304.7
Temporal Consistency. We compared Ca2-VDM with the two baselines (i.e., OS-Fix and OS-Ext) and existing SOTA autoregressive VDMs. To the best of our knowledge, existing autoregressive VDMs all use fixed-length conditional frames (similar to OS-Fix). We used Gen-L-Video (GenLV) [42] and StreamT2V [15]. Specifically, GenLV utilizes a base model AnimateDiff [12] and conducts co-denoising for overlapped 16-frame clips. We implemented it with an overlapping length (i.e., the condition length) of 8 frames. StreamT2V is based on Stable Video Diffusion (SVD) [2] and finetunes it conditioned on preceding frames to generate subsequent frames. It also generates 16 frames at each AR step, with 8 frames as the condition.
Figure 7: Qualitative examples from GenLV [42], StreamT2V [15], OS-Fix, and our Ca2-VDM. Yellow arrows highlight the consecutive frames having mutations.
@@ 4.2 Evaluation for Generation Quality
Figure 9: Results from OS-Ext and Ca2-VDM. They have comparable quality, while Ca2-VDM is more efficient in autoregressive generation, as evidenced in Table 10, Figure 8 and 11.
Figure 10: Cumulative time cost for autoregressively generating 80 frames at resolution 256$\times$256. Ext.C. means extendable condition. OS-Fix used $P$=8. OS-Ext and Ca2-VDM used $P_{\text{max}}$=25.
@@
We evaluated the FVD scores of each autoregression (AR) chunk w.r.t. the first chunk, as shown in Table 4. We can observe that Ca2-VDM has relatively lower FVD scores than the others. This indicates that extendable (long-term) condition helps to improve the temporal consistency. We also show qualitative examples in Figures 8. It shows content mutations in consecutive frames from the results of fixed-length condition methods, e.g., the 24th and 25th frames in GenLV, and the 65th and 66th frames in StreamT2V. We further compared Ca2-VDM with the condition extendable baseline, i.e., OS-Ext (cf. Figure 10). We see that Ca2-VDM shows comparable results with OS-Ext (while being more computationally efficient as demonstrated in the following Sec. 4.3).
@@ 4.3 Evaluation for Autoregression Efficiency
Time Cost. We first show the cumulative time cost of autoregressive generation in Table 10. Our models were tested on a single NVIDIA A100 GPU to generate 80 frames at resolution 256$\times$256, using improved DDPM [27] with 100 denoising steps. The result of StreamT2V [15] is from its GitHub page, which was tested on the same device and resolution. We can see that Ca2-VDM significantly improved over OS-Fix, OS-Ext, and StreamT2V [15], while being compatible with extendable condition. We further evaluated the accumulated time cost till each AR step, as shown in Figure 8. We can observe that: 1) Compared to OS-Fix, the time cost in Ca2-VDM has a clear reduction since it does not have redundant computations. 2) As the condition extends, the time cost of OS-Ext grows quadratically (before $P_{\text{max}}$ is reached), while the time cost of Ca2-VDM only grows linearly. 3) As the $P_{\text{max}}$ grows to incorporate longer condition, the increase of time cost for OS-Ext is significant, while it is relatively slight for Ca2-VDM.
@@ 5 Conclusions
Limitations are left in the Appendix.
@@
Table of Content
  • Sec. A: Limitations and Possible Future Directions
  • Sec. B: Training Objectives
@@
Appendix A Limitations and Possible Future Directions
Causal Modeling in Pretraining. Currently, all the pretrained weights for video diffusion models (either UNet-based, e.g., ModelScore-T2V [43], AnimateDiff [12], or Transformer-based, e.g., Open-Sora [55]) use bidirectional attention in their temporal modules. Our Ca2-VDM is built upon Open-Sora which was also pretrained using bidirectional attention. However, finetuning these bidirectionally pre-trained temporal modules using causal attention might be sub-optimal. The weights between bidirectional and causal temporal attention layers might have inherent gaps. Due to the limited computational resources, we did not conduct causal pretraining. Pretraining the VDM’s temporal modules from scratch (using causal attention) might have potential improvements.
Quality Degradation in Long-term Generation. As a common challenge, VDMs in long-term autoregressive generation suffer from frame appearance changes and quality degradation. Some works [15, 53] mitigate this issue by providing the VDM with the global appearance information extracted from the initial frame. However, during the long-term generation, video content may change and not all frames commit the same global appearance. In our setting, the long-term extendable context (i.e., early context from the KV-cache queue) helps mitigate the quality degradation, demonstrated by the results in Table 3 and Table 4. Further research on approaches addressing quality degradation is warranted and may hold potential significance for long-term video generation.
Appendix B Training Objectives
Recall that (cf. Sec. 3.2 in the main text) existing diffusion models [16, 27, 29] are trained with the variational lower bound of $\bm{z}_{0}$’s log-likelihood, formulated as
@@ Appendix B Training Objectives
Since $q$ and $p_{\theta}$ are both Gaussian, $D_{KL}$ is determined by the mean $\bm{\mu}_{\theta}$ and covariance $\bm{\Sigma}_{\theta}$. By re-parameterizing $\bm{\mu}_{\theta}$ as a noise prediction network $\bm{\epsilon}_{\theta}$ and fixing $\bm{\Sigma}_{\theta}$ as a constant variance schedule [16], the model can be trained using a simplified objective function:
@@
Following prior works [27, 29], we train the model with learnable covariance $\bm{\Sigma}_{\theta}$ to improve the sampling quality. This is achieved by optimizing the full $D_{KL}$ term in $\mathcal{L}_{\text{vlb}}$, resulting in an ~$\widetilde{\mathcal{L}}_{\text{vlb}}$ in our setting, i.e., applied with the same timestep vector $\bm{t}$ and loss mask $\bm{m}$. Then, the model is optimized by a combined loss function ~$\widetilde{\mathcal{L}}_{\text{simple}}+\widetilde{\mathcal{L}}_{\text{vlb}}$.

### v2

@@ 3.3 Autoregressive Inference with Cache Sharing
Figure 4: Illustration of causal temporal attention (a) & (b) and the temporal KV-cache queue with Cyclic-TPEs (c). In (c), $L_{\text{train}}=P_{\text{max}}+l$ and $P_{k+l}=P_{k}+l$. We show the state that autoregressive inference reaches $P_{k}=P_{\text{max}}$.
@@ 4.1 Experimental Setup
Model Details and Baselines. We built Ca2-VDM based on spatial-temporal Transformer following (Ma et al., 2025; Chen et al., 2024a) and initialized it with Open-Sora v1.0 (Zheng et al., 2024). Following PixArt-$\alpha$ (Chen et al., 2024a), we used T5 (Raffel et al., 2020) as the text encoder and used the VAE from StableDiffusion (Rombach et al., 2022). The length of the clean prefix was randomly sampled according to the multiples of chunk length $l$, i.e., $P\in\{1,1+l,\ldots,1+nl\}$ and $P_{\text{max}}=1+nl$. We used training videos of various lengths with $L_{\text{train}}=P+l$. As comparisons, we built two bidirectional baselines (cf. Figure 1(a)) based on the same Open-Sora v1.0: One was trained with fixed-length conditional frames (denoted as OS-Fix), where $P$ is fixed as $P=L_{\text{train}}/2$ in training and inference. The other was trained with autoregressively extendable conditional frames using the same training configs as Ca2-VDM (denoted as OS-Ext).
Table 1: Zero-shot FVD scores on MSR-VTT (Xu et al., 2016) and UCF101 (Soomro et al., 2012) test sets. All methods generate video at a resolution of 16$\times$256$\times$256. C: condition. T and I are text and image conditions, respectively.
@@
ModelScope (Wang et al., 2023b) T 550 410
VideoComposer (Wang et al., 2023c) T 580 -
Video-LDM (Blattmann et al., 2023b) T - 550.6
PYoCo (Ge et al., 2023) T - 355.2
Make-A-Video (Singer et al., 2023) T - 367.2
AnimateAnything (Dai et al., 2023) T+I 443 -
PixelDance (Zeng et al., 2024) T+I 381 242.8
SEINE (Chen et al., 2024b) T+I 181 -
@@
Table 2: Finetuned FVD scores on UCF-101 (Soomro et al., 2012) test set. Methods with were trained on both train and test sets.
@@
MCVD (Voleti et al., 2022) $64^{2}$ 1143
VDT (Lu et al., 2024) $64^{2}$ 225.7
DIGAN (Yu et al., 2022) $128^{2}$ 577
TATS (Ge et al., 2022) $128^{2}$ 420
VideoFusion (Luo et al., 2023) $128^{2}$ 220
LVDM (He et al., 2022) $256^{2}$ 372
PVDM (Yu et al., 2023) $256^{2}$ 343.6
Latte (Ma et al., 2025) $256^{2}$ 333.6
@@
[…]
Evaluation Datasets and Metrics. We used MSR-VTT (Xu et al., 2016), UCF101 (Soomro et al., 2012), and SkyTimelapse (Zhang et al., 2020) datasets at resolution 256$\times$256, and reported Fréchet Video Distance (FVD) (Unterthiner et al., 2019) following previous works (Zeng et al., 2024; Ge et al., 2023; Chen et al., 2024b). More details about choosing text prompts and computing FVD scores on these datasets are left Sec. D
4.2 Evaluation for Generation Quality
In-Chunk Generation Quality. We evaluated the zero-shot text-to-video (T2V) FVD scores on MSR-VTT (Xu et al., 2016) and UCF101 (Soomro et al., 2012), as shown in Table 2. We compared Ca2-VDM to state-of-the-art T2V models including two groups: 1) Text conditioned: ModelScope (Wang et al., 2023b), VideoComposer (Wang et al., 2023c), Video-LDM (Blattmann et al., 2023b), PYoCO (Ge et al., 2023), and Make-A-Video (Singer et al., 2023). 2) Text with extra image conditioned, e.g., for image-to-video: AnimateAnything (Dai et al., 2023), PixelDance (Zeng et al., 2024) and video transition: SEINE (Chen et al., 2024b). We also finetuned Ca2-VDM on UCF101 at resolution 16$\times$256$\times$256 and reported the FVD scores in Table 2. We compared it with SOTA video generation models: MCVD (Voleti et al., 2022), VDT (Lu et al., 2024), DIGAN (Yu et al., 2022), TATS (Ge et al., 2022), LVDM (He et al., 2022), PVDM (Yu et al., 2023), and Latte (Ma et al., 2025). The FVD results in both Table 2 and Table 2 show that our Ca2-VDM has a competitive T2V performance with SOTA models. More qualitative examples are left in Sec. E.
Table 3: FVD results on MSR-VTT test set.
Method FVD between AR step 1 and $i$
$i=2$ $i=3$ $i=4$ $i=5$ $i=6$
GenLV 282.8 291.4 299.0 318.2 310.3
StreamT2V 317.5 434.7 478.2 462.0 512.4
OS-Fix 182.9 210.6 260.8 284.3 315.1
Ca2-VDM 160.6 206.5 262.8 281.3 304.7
Temporal Consistency. We compared Ca2-VDM with the two baselines (i.e., OS-Fix and OS-Ext) and existing SOTA autoregressive VDMs. To the best of our knowledge, existing autoregressive VDMs all use fixed-length conditional frames (similar to OS-Fix). We used Gen-L-Video (GenLV) (Wang et al., 2023a) and StreamT2V (Henschel et al., 2025). Specifically, GenLV utilizes a base model AnimateDiff (Guo et al., 2024b) and conducts co-denoising for overlapped 16-frame clips. We implemented it with an overlapping length (i.e., the condition length) of 8 frames. StreamT2V is based on Stable Video Diffusion (Blattmann et al., 2023a) and finetunes it conditioned on preceding frames to generate subsequent frames. It also generates 16 frames at each AR step, with 8 frames as the condition.
Table 4: Ablations of $P_{\text{max}}$ and prefix-enhancement (PE) on SkyTimelapse (Zhang et al., 2020). Each variant of Ca2-VDM generated 48 frames by 6 AR steps. The results were divided into three 16-frame chunks for FVD evaluation.
@@ 4.2 Evaluation for Generation Quality
Figure 5: Qualitative examples from GenLV (Wang et al., 2023a), StreamT2V (Henschel et al., 2025), OS-Fix (Zheng et al., 2024), and Ca2-VDM. Yellow arrows highlight consecutive frames having mutations.
@@
Figure 7: Results from OS-Ext and Ca2-VDM. They have comparable quality, while Ca2-VDM is more computationally efficient, as evidenced in Table 8, Figure 6 and 9.
Figure 8: Time cost for generating 80 frames at resolution 256$\times$256. OS-Fix used $P$=8. OS-Ext and Ca2-VDM used $P_{\text{max}}$=25. Ext.C. means extendable condition.
@@
We evaluated the FVD scores of each autoregression (AR) chunk w.r.t. the first chunk, as shown in Table 3. We can observe that Ca2-VDM has relatively lower FVD scores than the others. This indicates that extendable (long-term) condition helps to improve the temporal consistency. We also show qualitative examples in Figures 6. It shows content mutations in consecutive frames from the results of fixed-length condition methods, e.g., the 24th and 25th frames in GenLV, and the 65th and 66th frames in StreamT2V. We further compared Ca2-VDM with the condition extendable baseline, i.e., OS-Ext (cf. Figure 8). We see that Ca2-VDM shows comparable results with OS-Ext (while being more computationally efficient as demonstrated in Sec. 4.3). We conducted further comparisons between Ca2-VDM and OS-Ext in terms of video quality and long-term content drift. The results are left in Sec. E of the Appendix.
@@ 4.3 Evaluation for Autoregression Efficiency
Time Cost. We first show the cumulative time cost of autoregressive generation in Table 8. Our models were tested on a single NVIDIA A100 GPU to generate 80 frames at resolution 256$\times$256, using improved DDPM (Nichol & Dhariwal, 2021) with 100 denoising steps. The result of StreamT2V (Henschel et al., 2025) is from its GitHub page, which was tested on the same device and resolution. We can see that Ca2-VDM significantly improved over OS-Fix, OS-Ext, and StreamT2V (Henschel et al., 2025), while being compatible with extendable condition. We further evaluated the accumulated time cost till each AR step, as shown in Figure 6. We can observe that: 1) Compared to OS-Fix, the time cost in Ca2-VDM has a clear reduction since it does not have redundant computations. 2) As the condition extends, the time cost of OS-Ext grows quadratically (before $P_{\text{max}}$ is reached), while the time cost of Ca2-VDM only grows linearly. 3) As the $P_{\text{max}}$ grows to incorporate longer condition, the increase of time cost for OS-Ext is significant, while it is relatively slight for Ca2-VDM.
@@
Table 5: GPU memory usage comparison between Live2diff (Xing et al., 2024) and Ca2-VDM. The comparisons are not strictly aligned since Live2diff is Unet-based. The resolution of the generated video is $256\times 256$. $L$ is the number of generated frames at each auto-regression step. $H$ and $W$ are after 8×8$\times$8 × VAE down sampling. The values of $h^{\prime}w^{\prime}$ and $C^{\prime}$ vary across blocks due to the down-sampling and up-sampling in Unet. PE means prefix-enhancement (cf.  Eq.(4)).
Method Denoising Model Forward Shape KV-cache Shape KV-cache Total
Steps ($T$) ($B,C,L,H,W$) ($T,L_{\text{cond}},hw,C^{\prime}$) Memory Cost Memory Cost
Live2diff 4 (4 , 4, 1, 32, 32) (4, 16, $h^{\prime}w^{\prime}$, $C^{\prime}$) 1.42 GB 10.90 GB
Live2diff 50 (50, 4, 1, 32, 32) (50, 16, $h^{\prime}w^{\prime}$, $C^{\prime}$) 17.70 GB 29.46 GB
Ca2-VDM w/ PE 50 (1 , 4, 8, 32, 32) (1, 25, $hw$, $C$) 0.86 GB 4.79 GB
Ca2-VDM w/o PE 50 (1 , 4, 8, 32, 32) (1, 25, $hw$, $C$) 0.77 GB 3.95 GB
Memory Cost. We conducted empirical GPU memory statistics, as shown in Table 5. We compared Ca2-VDM with a concurrent work, Live2diff (Xing et al., 2024). It stores KV-cache for every denoising step (with different noise levels $t$ and thus different KV features), which costs much more GPU memory than ours. Note that Live2diff uses a batch size that is equal to the number of denoising steps, i.e., $B=T$. This is because it uses pipeline denoising following StreamDiffusion (Kodaira et al., 2023), which puts frames with progressive noisy levels into a batch and generates one frame each autoregression step. Benefited from cache sharing, Ca2-VDM’s memory cost is independent of denoising steps, as its fixed shape $(1,25,hw,C)$ ensures constant memory usage. In contrast, Live2diff’s memory cost scales with $T$ (e.g., from 1.42 GB at $T=4$ to 17.70 GB at $T=50$), confirming that cache sharing saves $T\times$ GPU memory. As a result, Ca2-VDM requires only 0.86 GB (w/ PE) or 0.77 GB (w/o PE), with the difference due to spatial KV-cache for prefix-enhancement (PE).
@@ 5 Conclusions
Acknowledgements
This work was supported by the National Key Research & Development Project of China (2024YFB3312900), Key R&D Program of Zhejiang (2025C01128), an Fundamental Research Funds for the Central Universities. Long Chen was supported by the Hong Kong SAR RGC Early Career Scheme (26208924), the National Natural Science Foundation of China Young Scholar Fund (62402408), Huawei Gift Fund, and the HKUST Sports Science and Technology Research Grant (SSTRG24EG04). Kaifeng Gao was supported by the 2024-2025 Grant for Pursuing Outstanding Doctoral Dissertations of Zhejiang University.
Impact Statement
Our Ca2-VDM is a generic fast video generation paradigm. It is potentially powerful to boost existing VDMs to generate high-quality live-stream videos. The live-stream (or real-time) video generation techniques have a revolutionary impact on the field of content creation industry, and have great potential commercial values. Meanwhile, it’s necessary to note that Ca2-VDM also has the inherent risks of common image/video generation models, such as generating videos with harmful or offensive content, or being used by malicious actors for generating fake news. We can use some watermarking technologies (e.g., (Lukas & Kerschbaum, 2023)) to avoid the generated videos being abused.
@@ Acknowledgements
  • Sec. A: Illustration of Prefix-enhanced Spatial Attention
  • Sec. B: Detailed Training Objectives
@@
  • Sec. E: More Experiment Results
  • Sec. F: Limitations and Possible Future Directions
Appendix A Illustration of Prefix-enhanced Spatial Attention
We provide more details of Prefix-enhanced Spatial Attention (cf.  Eq. (4)) in Figure 10.
Appendix B Detailed Training Objectives
Recall that (cf. Sec. 3.2 in the main text) existing diffusion models (Ho et al., 2020; Nichol & Dhariwal, 2021; Peebles & Xie, 2023) are trained with the variational lower bound of $\bm{z}_{0}$’s log-likelihood, formulated as
@@ Appendix B Detailed Training Objectives
Since $q$ and $p_{\theta}$ are both Gaussian, $D_{KL}$ is determined by the mean $\bm{\mu}_{\theta}$ and covariance $\bm{\Sigma}_{\theta}$. By re-parameterizing $\bm{\mu}_{\theta}$ as a noise prediction network $\bm{\epsilon}_{\theta}$ and fixing $\bm{\Sigma}_{\theta}$ as a constant variance schedule (Ho et al., 2020), the model can be trained using a simplified objective function:
@@
Following prior works (Nichol & Dhariwal, 2021; Peebles & Xie, 2023), we train the model with learnable covariance $\bm{\Sigma}_{\theta}$ to improve the sampling quality. This is achieved by optimizing the full $D_{KL}$ term in $\mathcal{L}_{\text{vlb}}$, resulting in an ~$\widetilde{\mathcal{L}}_{\text{vlb}}$ in our setting, i.e., applied with the same timestep vector $\bm{t}$ and loss mask $\bm{m}$. Then, the model is optimized by a combined loss function ~$\widetilde{\mathcal{L}}_{\text{simple}}+\widetilde{\mathcal{L}}_{\text{vlb}}$.